OpenStudy (anonymous):

1/2 / 1/4 = ???? please help

OpenStudy (anonymous):

isn't it 2?

OpenStudy (anonymous):

can anybody help here?

OpenStudy (anonymous):

is the answer 2?

OpenStudy (anonymous):

Its 2 heres the steps to getting the answer, http://www.cymath.com/answer.php?q=.50%2F.25

OpenStudy (anonymous):

2/8

OpenStudy (anonymous):

I could most likely tell you the answer but I can't explain it. I'll let someone else help .:)

OpenStudy (anonymous):

@e.mccormick @karatechopper i did not ask for direct answer.

OpenStudy (anonymous):

@karatechopper they given me direct answers. ;(

OpenStudy (anonymous):

HAHA! I was catching cheaters so I could get a mod on here to take care of you! :3

OpenStudy (karatechopper):

Then ask for them to explain their answer...

OpenStudy (karatechopper):

Jokes on you darling.

OpenStudy (anonymous):

That is not right, that is not what the code of conduct says.

OpenStudy (anonymous):

mod don't know what code of conduct says? http://openstudy.com/code-of-conduct read it

OpenStudy (anonymous):

Don’t post only answers - guide the asker to a solution.

OpenStudy (anonymous):

As long as they explain their answer, I don't think it should be against the rules.

OpenStudy (anonymous):

Don't be offensive, inappropriate, or creepy! So avoid: Insulting, berating, or otherwise discouraging others.

OpenStudy (anonymous):

you are lacking a brain to read the code of conduct.

OpenStudy (anonymous):

Don't provide someone with just the answer - explain the process, and help guide them through understanding the problem. Don’t post ONLY answers - guide the asker to a solution. Have homework problems? We can help you - as long as you are looking to learn how to get your solution (and not just the answer). I don't see anything wrong here. He didn't just give you the answer to the problem. He also gave you a link to explain how he got it.

OpenStudy (anonymous):

i thought it was openstudy not opencheating.

OpenStudy (anonymous):

you gave a direct answer just to get a medal not because you like helping, that is called cheating.

OpenStudy (anonymous):

but i don't medal cheaters

OpenStudy (anonymous):

openstudy needs less cheaters and more helpers.

OpenStudy (anonymous):

you are the one who is going around giving away direct answers just for medals. speak for yourself, you're going to let him do this? @karatechopper I thought you were a moderator, why are you agreeing with them?

OpenStudy (anonymous):

No one here is cheating. You asked the question. He didn't only give you the answer. He also gave you an explanation. Did you look at the link. Yes, they did use a calculator to solve it but it ALSO explains it. @a.prime did NOTHING wrong. The only one who did something wrong is you if you took the answer without knowing how to do it and ask for more of an explanation. @Solver arguing about it isn't going to solve anything. We have moderators for a reason. @Solver you should let them do their job.

OpenStudy (anonymous):

I didn't know openstudy would choose someone who agrees with cheaters to become a moderator.

OpenStudy (anonymous):

The moderator isn't doing his/her job and I was using the question as bait to catch cheaters, I didn't look at the link and if he did use a calculator he didn't not put any effort into 'helping' me. 'Explanation' doesn't mean anything to me neither should it mean anything to ANYONE on openstudy.

OpenStudy (haleyelizabeth2017):

Okay solver, it's okay........just ignore this now and try to move on....:/ arguing with them won't change their opinions......and it won't get you anywhere.

OpenStudy (anonymous):

But I do no 'GUIDING THE ASKER INTO THE SOLUTION TO THE PROBLEM' WHICH is stated in the CODE OF CONDUCT.

OpenStudy (anonymous):

You're right. @haleyelizabeth2017 A cheater is a cheater.

OpenStudy (anonymous):

I can't do anything about that. Sorry, but he didn't know how to help I guess. I don't know... We should just ignre it and when you see someone do it again just report them. :)

OpenStudy (haleyelizabeth2017):

please just stop

OpenStudy (anonymous):

That's all we can do.

OpenStudy (anonymous):

Reporting won't do them good if mods agree with them. >> Question Closed <<

OpenStudy (karatechopper):

Ok. calm down.

OpenStudy (karatechopper):

Using this as an incentive to get others in trouble is also against the code of conduct. Be nice. I've taken my action that is part of my responsibility, now you may take yours as a user.

OpenStudy (karatechopper):

No need to argue.

OpenStudy (karatechopper):

And once again, I have already taken action.

OpenStudy (anonymous):

No no no, I'M HELPING YOU DO YOUR JOB. CATCHING CHEATERS, YOU'RE TELLING ME IT'S GOOD TO LET CHEATERS GET BY WITH THING ON HERE?

OpenStudy (anonymous):

@phi

OpenStudy (phi):

1/2 / 1/4 \[ \frac{ \frac{1}{2}}{\frac{1}{4} } \] The "fast way" is to remember this rule: instead of dividing by a number (here 1/4), "flip the number and multiply" in other words rather than divide by 1/4 we "flip it" to get 4/1 which simplifies to just 4 now multiply 1/2 * 4 = 2 another way to do it is multiply the top and bottom by the same number , 4 in this case \[ \frac{ \frac{1}{2}\cdot 4}{\frac{1}{4} \cdot 4} \] notice that the bottom is 1/4 * 4 which is 1 : \[ \frac{ \frac{1}{2}\cdot 4}{1} \] when we divide by 1 we can simplify to \[ \frac{1}{2}\cdot 4 \] and now do 1/2 * 4 \[ \frac{1}{2}\cdot 4 =\frac{1 \cdot 4}{2} = \frac{4}{2}=2\]

OpenStudy (karatechopper):

I'm not saying its bad, but if you are actually catching the cheaters, then I wouldn't really call this cheating. They gave you an answer, yes I understand that you reported them for giving a direct answer, but you can also ask "how did you get that". I see your point, I just don't believe this sort of way to catch people off guard is the best way.

OpenStudy (anonymous):

Agreed.

OpenStudy (anonymous):

@phi Haven't you been reading this whole conversation....

OpenStudy (anonymous):

Agree